Starling Co. is considering disposing of a machine with a book value of $21,500 and estimated remaining life of five years. The old machine can be sold for $5,300. A new high-speed machine can be purchased at a cost of 69,300. It will have a useful life of five years and no residual value. It is estimated that the annual variable manufacturing costs will be reduced from $22,600 to $19,500 if the new machine is purchased. The differential effect on income for the new machine for the entire five years is

a.increase of $63,050

b.decrease of $48,500

c.increase of $48,500

d.decrease of $63,050
Answer #1

Differential effect on income :

Old machine New machine
Salvage value -5300
Purchase cost 69300
Variable manufacturing cost 22600*5 = 113000 19500*5 = 97500
Total relevant cost 113000 161500

Operating income decrease by = 113000-161500 = -48500

So answer is b) Decrease by $48500